OpenStudy (anonymous):

Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (anonymous):

$0.50 to purchase sale price = .50(1.40) = $ 0.70 this means they make $ 0.20 profit per flower 0.20f = 500 f = 500/0.20 f = 2500 Angelica's service club would have to sell 2500 flowers to clear a profit of $500.

OpenStudy (anonymous):

Thank you so much. I honestly suck at algebra

OpenStudy (anonymous):

any more questions

OpenStudy (anonymous):

yes. Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. 3x+6 8 = 7x−1 6 24 1 ( 3x+6 8 ) = 24 1 ( 7x−1 6 ) 3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 = 28x−9x−1 18 = 19x−1 18+1 = 19x−1+1 19 = 19x 1 = x

OpenStudy (anonymous):

can you give me medal

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Is the original problem: (3x + 6) / 8 = (7x - 1) / 6 To check to see if x = 1 is correct, substitute the "1" for the "x" in the original problem and see if it works out: (3(1) + 6) / 8 = (7(1) - 1) / 6 (3 + 6) / 8 = (7 - 1) / 6 9 / 8 = 6 / 6 9 / 8 = 1 FALSE So x = 1 is not the solution. And here is where the problem is: 3(3x+6) = 4(7x−1) 9x+18 = 28x−1 <--- You multiplied 4(7x - 1) to get 28x - 1, where it should have been 28x - 4.

OpenStudy (anonymous):

thank yooou. and then this one Angelo took a group of people to the movies and spent $80 on tickets. Adult tickets cost $10, and child tickets cost $6. The equation to represent his cost is 10A + 6C = $80. If there were 5 children in the group, how many adult tickets did he purchase? Show your work and use a complete sentence to explain your solution in the context of the original problem.

OpenStudy (ryanwhite499):

Make sure you post this in math. A lot of people well help you more than we can

OpenStudy (anonymous):

@ryanwhite499 thanks sorry

OpenStudy (anonymous):

5 adults 10a+6(5)=80 10a+30=80 10a-30+30=80-30 10a=50 10a/10=50/10 a=5 there are six children, 30 must be spent, 50 is left, since each adult costs 10, 5 there are 5 adult tickets

OpenStudy (anonymous):

any more and if there is go to the math section

OpenStudy (anonymous):

do you know how

OpenStudy (anonymous):

yes. just one more, though.

OpenStudy (anonymous):

i will help you there

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wat is it

OpenStudy (anonymous):

A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (anonymous):

515 is the minimum diamter 535 is the maximum its just addition and subtraction, 525+10, 525-10, please don't forget the decimal points.

OpenStudy (anonymous):

where are the decimal points?

OpenStudy (anonymous):

| x - 0.525| = 0.01 x - 0.525 = 0.01 x = 0.01 + 0.525 x = 0.535 -| x - 0.525| = 0.01 -x + 0.525 = 0.01 -x = 0.01 - 0.525 -x = 0.515 x = - 0.515 minimum is : -0.515 diameter maximum is : 0.535 diameter

OpenStudy (anonymous):

oh okaaaaay i get it

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

np